\addcontentsline{toc}{chapter}{Bibliografia}
\begin{thebibliography}{10}
       \bibitem{Beck} J. Beck. \emph{Diophantine approximation and quadratic fields}. Number Theory, K. Gyory, A. Petho and V. T. Sós (eds.), de Gruyter, 1998, 55-93.
       \bibitem{Biro1} A. Biró. \emph{Yokoi's Conjecture}. Acta Arithmetica 106 (2003), no. 1, 85–104.
       \bibitem{Biro3} A. Biró. \emph{Chowla's Conjecture}. Acta Arithmetica 107 (2003), no. 2, 179–194.
       \bibitem{Biro2} A. Biró. \emph{Class number one problem for a family of real quadratic fields}. Mathematika 68 (2022), no. 4, 1221–1232.
       \bibitem{BiroGranville} A. Biró, A. Granville. \emph{Zeta functions for ideal classes in real quadratic fields, at s=0}. J. Number Theory 132 (2012), no. 8, 1807-1829.
       \bibitem{Cherubini} G. Cherubini, A. Fazzari, A. Granville, V. Kala, and P. Yatsyna. \emph{Consecutive Real Quadratic Fields with Large Class Numbers}. International Mathematics Research Notices IMRN 2023, no. 14, 12052–12063.
       \bibitem{Davenport} H. Davenport. \emph{Multiplicative number theory}, vol. 74, Springer Sciences \& Business Media, Berlin, 2013.
       \bibitem{Goldfeld} D. Goldfeld. \emph{Gauss' class number problem for imaginary quadratic fields}. Bulletin (New Series) of the
American Mathematical Society, Volume 13, Number 1, July 1985.
       \bibitem{H-K 1} Von F. Halter-Koch. \emph{Einige periodische Kettenbruchentwicklungen und Grundeinheiten quadratischer Ordnungen}. Abhandlungen aus dem Mathematischen Seminar der Universität Hamburg
 59, 157-169 (1989).
       \bibitem{Lamzouri} Y. Lamzouri. \emph{Extreme values of class numbers of real quadratic fields}. International Mathematics Research Notices IMRN (2015), no. 22, 11847–11860.
       \bibitem{McZimmer} J. Mc Laughlin, P. Zimmer. \emph{Some more long continued fractions, I}. Acta Arithmetica 127 (2007), no. 4, 365–389.
       \bibitem{Schmidt} W. M. Schmidt. \emph{Approximation to irrational numbers by rationals}, Diophantine Approximation. Springer, Berlin, 1980, pp. 1-26.
\end{thebibliography}

\documentclass[a4paper,10pt]{amsart}

\usepackage[english]{babel}
\usepackage[utf8]{inputenx}
\usepackage[T1]{fontenc} 
\usepackage{graphicx}         
\usepackage{xcolor}  
\usepackage{hyperref}
\usepackage{tikz-cd}
\usepackage{lmodern}
\usepackage{verbatim}
\usepackage{longtable}

\title{Class number of real quadratic fields of explicit discriminant}

\author{Riccardo Bernardini}
\address{Department of Mathematics, University of Rome ``La Sapienza'', Rome, Italy}
\email{r.bernardini@uniroma1.it}
\date{\today}  

\subjclass[2020]{Primary 11R29; Secondary 11R11, 11A55}  
\keywords{Real quadratic fields, class number, continued fractions}  

\makeatletter
\newcommand{\addresshere}{%
  \enddoc@text\let\enddoc@text\relax
}
\makeatother

\begin{document}
\newtheorem{thm}{Theorem}[section]
\newtheorem{defn}[thm]{Definition}
\newtheorem{lem}[thm]{Lemma}
\newtheorem{cor}[thm]{Corollary}
\newtheorem{prop}[thm]{Proposition}
\newtheorem{ex}[thm]{Example}
\theoremstyle{definition}\newtheorem{rmk}[thm]{Remark}

\begin{abstract}
In this paper we are interested in the class numbers of a family of real quadratic fields for which the square roots of the discriminants have a known expansion in continued fraction. In particular we prove that $h(D)>1$, with possibly a finite number of exceptions.
\end{abstract}

\maketitle

\section{Introduction}
One of the most important invariants of a number field is its class number. This is quite mysterious and, in spite of many efforts and abundance of literature, there are still lots of unsolved problems. For instance, Gauss' famous conjecture asserts the existence of infinitely many real quadratic fields with class number one. 
Gauss' Conjecture is notably difficult to attack: as shown by the class number formula, the essence of the issue is to find good estimates for the growth rate of the fundamental unit in relation to the discriminant. Indeed, if $\mathbb{Q}(\sqrt{D})$ has discriminant $D$ and class number $h(D)$, the class number formula and an ineffective theorem by Siegel (see Davenport \cite{Davenport}) provide the following inequality
\begin{equation}\label{diseq an 1}
h(D)\gg_\epsilon \frac{D^{\frac{1}{2}-\epsilon}}{\log{\xi_D}},
\end{equation}
where $\xi_D$ is the fundamental unit and $\gg_\epsilon$ means that the left-hand side is greater than the right-hand side multiplied by a positive constant depending on $\epsilon$. 
Clearly if $\xi_D$ does not grow too fast in relation to $D$ the class number will be eventually bigger than any constant. We have plenty of examples of this: the are many explicit families with $\log{\xi_D}$ growing like $\log{D}$ or $(\log{D})^2$, so the class number tends to $\infty$. Thus two interesting questions arise. One is: what is the greatest order of growth of $h(D)$? And the other is: for the explicit families of discriminants available in the literature (see for example Halter-Koch \cite{H-K 1} or McLaughlin and Zimmer \cite{McZimmer}) from what point on do they produce class numbers greater than one?
Regarding the first question we mention Lamzouri's \cite[Theorem 1.2]{Lamzouri}, where he gives a very precise quantitative result on class numbers as large as possible, i.e. he proved that
\begin{equation}\label{ineq Cher}
h(D)\geq (2e^\gamma + o(1))\sqrt{D}\frac{\log{\log{D}}}{\log{D}},
\end{equation}
where $\gamma$ is the Euler-Mascheroni constant, is verified by at least $x^{1/2-1/\log\log{x}}$ and at most $x^{1/2+o(1)}$ fundamental discriminants $\leq x$. Some years later, in 2023, it was proved in \cite{Cherubini} that one can produce an arbitrary number of consecutive real quadratic fields with class number satisfying an inequality like \eqref{ineq Cher}.

In this paper we focus on the second question and we consider the family of real quadratic fields with discriminants described in \cite[Theorem 6]{McZimmer} by McLaughlin and Zimmer
\begin{align}\label{D McZim}
D(b,s,k)=(4bs+1)^k+(b(4bs+1)^k+s)^2
\end{align}
with $b,s,k>0$. 
Our main result is

\begin{thm}\label{thm 1}
Let $b\geq0$ and $s,k>0$ be such that $b+s$ is odd and $D=D(b,s,k)$ (as in equation \eqref{D McZim}) is squarefree. If $41\cdot61\cdot175\cdot1861|b$, then $h(D)>1$ unless $b=0$ and $s=1$.
\end{thm}
We immediately explain the reasons behind the choice of $D(b,s,k)$ and the numbers $41,61,175,1861$ in the statement of Theorem \ref{thm 1}. 
As the reader will notice, Theorem \ref{thm 1} is in the same fashion of \cite[Theorem 1.1]{Biro2} where Biró considered the family
\begin{align}\label{D Biro}
D_{Biró}(b,s,k)=4(bs+1)^k + (b(bs+1)^k+s)^2
\end{align}
with $b\geq0,s>0,k\geq 2$ under the hypothesis that it is squarefree (this also implies $D\equiv 1 \pmod{4}$).
One common aspect between the families \eqref{D McZim} and \eqref{D Biro} is that, if $(1,\omega_D)$ is the standard integral basis of the ring of integers of $\mathbb{Q}(\sqrt{D})$ with $D=D(b,s,k),D_{Biró}(b,s,k)$, i.e. $\omega_D=(1+\sqrt{D})/2$ if $D\equiv 1 \pmod{4}$ and $\sqrt{D}$ otherwise, and
\begin{equation*}
\omega_D=[b_0,\overline{b_1,\ldots,b_l}],
\end{equation*}
the coefficients $b_j$ for $j\neq 0,l$ are divisible by $b$. A slight difference is that the continued fraction of the square root of \eqref{D McZim} has a factor 2 which appears in the coefficients; we will meet this factor again throughout the proofs. 
About the numbers $41,61,175,1861$, they are so chosen because they behave well in relation to the computations. As it will be better explained in section \ref{outline} the strategy is to progressively erase residue classes by mean of explicit Dirichlet characters of conductors $41,61,175,1861$; since the choice of these characters is not unique, it is possible to exhibit different finite sets of numbers which make Theorem \ref{thm 1} (and also \cite[Theorem 1.1]{Biro2}) work. 
Furthermore, Biró got information about \eqref{D Biro} by looking at it  modulo a prime $q$ that divides $b$ and so reducing to the case of Yokoi's family $n^2+4$, which he had previously studied in \cite{Biro1}. On the other hand, \eqref{D McZim} modulo a prime $q$ dividing $b$ becomes $n^2+1$. 
This suggests to analyze firstly the discriminants of the form $n^2+1$ and then to deduce information about \eqref{D McZim} as to Biró did in his papers.
Another analogy is that both $\mathbb{Q}(\sqrt{n^2+4})$ and $\mathbb{Q}(\sqrt{n^2+1})$, with odd $n$, have fundamental unit with negative conjugate: we will need this in a technical step that involves Shintani's zeta function into the computations (see section \ref{family nn+1}).
The condition $b+s\equiv 1 \pmod{2}$ is required so that $D\equiv 2 \pmod{4}$: if $D(b,s,k)\equiv 1 \pmod{4}$ we need the expansion of $(1+\sqrt{D(b,s,k)})/2$ which leads to a bit harder computations and we loose some analogies with the family \eqref{D Biro}.

We put emphasis on the similarities between \eqref{D McZim} and \eqref{D Biro} because it is important to fulfill some requirements in order to use Biró's strategy at full strength. In fact, in spite of being quite ingenious, his methods seem to be built \textit{ad hoc}.

We end the introduction with some final remarks and results.
\begin{rmk}
If $bsk=0$ then the discriminant $D(b,s,k)$ in \eqref{D McZim} becomes $n^2+1$ where $n=b+s$. If $n$ is even it has already been investigated by Biró in \cite{Biro3}, who exhibited the complete list of real quadratic fields with class number one. If $n$ is odd we will give a proof in section \ref{family nn+1} that the class number is always at least two with one exception (see Corollary \ref{cor 2 thm 2}).
\end{rmk}

We also notice that if $(b,s)>2$ with $b,s>0$ or $4bs+1$ is not a prime, then Theorem \ref{thm 1} can be drastically improved by removing almost all the hypotheses and giving a complete characterization of the values of $D$ for which $h(D)=1$; this is the statement of the next two propositions, which will be proved in section \ref{proof prop 1}.

\begin{prop}\label{prop 1}
If $D=D(b,s,k)$ squarefree is as in equation \eqref{D McZim} and $(b,s)>2$, then $h(D)>1$.
\end{prop}

\begin{prop}\label{prop 2}
If $D=D(b,s,k)$ squarefree is as in equation \eqref{D McZim} and $4bs+1$ is not a prime, then $h(D)>1$.
\end{prop}
Here we exclusively rely on the theory of continued fractions and we use a method, already present in \cite{Biro2}, that we apply also in the end of section \ref{proof thm 1}.

Since Biró in \cite{Biro2} uses some computations already done in \cite{Biro1}, in section \ref{family nn+1} we perform the analogous ones by PARI/GP for the family $n^2+1$ and then we use them in section \ref{proof thm 1} to prove Theorem \ref{thm 1}. An example of the code will be shown in the appendices. 

We report the main result about the family $n^2+1$ which corresponds to the theorem shown in \cite[§1]{Biro1}. 

\begin{thm}\label{thm 2}
If $D=n^2+1$, with $n$ odd, is squarefree and $h(D)=1$, then $D$ is a quadratic residue modulo at least one of the following: $5,7,41,61,1861$.
\end{thm}

The fact that the prime factors of the numbers appearing in Theorem \ref{thm 1} and Theorem \ref{thm 2} are the same is not a coincidence: indeed, as we pointed out above, if $q|b$, then $D(b,s,k)\equiv s^2+1 \mod{q}$.
A consequence of Theorem \ref{thm 2} is that, for $n$ odd and greater than an explicit bound, $h(D)>1$; this is stated in the next corollary.

\begin{cor}\label{cor 1 thm 2}
If $D=n^2+1$, with $n$ odd, is squarefree and $931=\lceil\frac{1861}{2}\rceil\leq n$ then $h(D)>1$.
\end{cor}
A computational check with PARI/GP completes the study of the family $n^2+1$ giving the full classification.

\begin{cor}\label{cor 2 thm 2}
If $D=n^2+1$, with $n$ odd, is squarefree, then $h(D)>1$ if and only if $n>1$.
\end{cor}
By Corollary \ref{cor 2 thm 2} and by the paragraph following the Corollary of  \cite[§1]{Biro3} one gets all the values of $n$ for which $n^2+1$ is squarefree and $h(n^2+1)=1$.
\begin{cor}
If $n^2+1$ is squarefree then $h(n^2+1)=1\iff n=1,2,4,6,10,14,26$.
\end{cor}

\section{Outline of the proof}\label{outline}

Throughout the rest of the article we will make use of the following notation:
$K$ is the real quadratic field $\mathbb{Q}(\sqrt{D})$ with $D$ squarefree, $d$ is its discriminant, $\mathcal{O}_K$ is the maximal order (i.e. the ring of integers) of $K$, $\xi_D$ (or simply $\xi$ when there is no confusion) is the fundamental unit of $K$, $\chi_d$ is the Kronecker character and $h(D)$ (or simply $h$ when there is no confusion) is the class number. By $\zeta(s)$ we mean the Riemann zeta function, by $\zeta_K(s)$ the Dedekind zeta function, while if $\chi$ is a Dirichlet character we define the twisted Dedekind zeta function as
$$\zeta_K(s,\chi):=\sum_{0\neq\mathfrak{a}\subset\mathcal{O}_K}\frac{\chi(N(a))}{N(a)^s}.$$
Similarly, if we indicate by $P(K)$ the set of principal integral ideals, we can define
\begin{equation*}
\zeta_{P(K)}(s):=\sum_{\frak{a}\in P(K)}\frac{1}{N(a)^s}\quad \mbox{and}\quad \zeta_{P(K)}(s,\chi):=\sum_{\frak{a}\in P(K)}\frac{\chi(N(a))}{N(a)^s}
\end{equation*}
Furthermore, for a given character $\chi$, by $\mathbb{Q}(\chi)$ we mean the number field obtained by adding the values $\chi(a)$ to $\mathbb{Q}$ for every integer $a$.

The first family we are going to study is $n^2+1$ with odd $n$ so that it produces numbers equivalent to $2$ modulo $4$. Thus, if they are squarefree, the standard integral basis is $(1,\sqrt{D})$ and the discriminant of the field $d=4D$. If $\chi$ is a Dirichlet character modulo $q$ with some further properties we construct two algebraic integers as in \cite{Biro1}, namely 
\begin{equation}\label{def m}
m_\chi:=\sum_{a=1}^qa\chi(a)
\end{equation}
and
\begin{equation}\label{def A}
A_\chi(n):=\sum_{0\leq C,D\leq q-1}\chi(D^2-C^2-nCD)\left\lceil\frac{nC-D}{q}\right\rceil\left(C-q\right).
\end{equation}
This crucial step, stated in Lemma \ref{lemma fatto A}, is proved through identities involving the $L$-functions and the zeta functions.
Now suppose $n$ is such that $n^2+1$ is not a square residue for a prime $p$ dividing $q$. Assume moreover that there exists a prime ideal $I$ of $\mathbb{Q}(\chi)$ containing $m_\chi$ laying above an integer prime $r$ different from $q$ and $2$. All this forces $n$ to lay in a precise residue class modulo $I$,
indeed we have that $A_\chi(2n)\in I$ by Lemma \ref{lemma fatto A} and, if we write $n=Pq+n_0$ by the Euclidean division, we get
$$
0\equiv A_\chi(2n)=A_\chi(2n_0)+2PB_\chi(2n_0)\quad (I),
$$ 
where 
\begin{equation}\label{def B}
B_\chi(a):=\sum_{0\leq C,D\leq q-1}\chi(D^2-C^2-aCD)C(C-q),
\end{equation}
and so 
$$
n\equiv n_0 -q\frac{A_\chi(2n_0)}{2B_\chi(2n_0)}\quad (I).
$$
If we choose $I$ such that $\mathcal{O}_{\mathbb{Q}(\chi)}/I\simeq \mathbb{Z}/(r)$, $n$ is then forced to be in a certain class modulo $r$, say $r_0$. We can know ask ourselves if $r_0^2+1$ is a square modulo $r$ or not. If not we find an ideal $I'$ over another prime $r'$ finding new constraints for $n$. This process continues until we find a finite set of numbers so that $n^2+1$ is a square modulo at least one of them. With the aim of making the paper more readable we put the computations in the appendices in the end of the paper.

\begin{rmk}\label{rmk 2.1}
As in \cite{Biro1} the reason we can rely on computation, of which we will insert an example of the PARI/GP program in appendix \ref{app B}, is that we can choose $I$ so that $\mathcal{O}_{\mathbb{Q}(\chi)}/I\simeq \mathbb{Z}/(r)$. Thus we are actually working with integers.
\end{rmk}

We point out that the idea of excluding residue classes, which is behind this article and the works of Biró \cite{Biro1}, \cite{Biro2} and \cite{Biro3} go back until Beck's paper \cite{Beck}. Here he combined formulas for special values of zeta functions associated to the number field or suitable quadratic characters and he could exclude some residue classes. Biró in \cite{Biro1} understood we could enhance his method by involving zeta functions related to nonquadratic Dirichlet characters, so that he had a greater amount of possibilities for erasing residue classes.

The next part of the paper, section \ref{proof thm 1}, begins with a series of technical results consisting in an application to our case of the work of Biró and Granville \cite{BiroGranville}. The general idea is to make use again of the algebraic integer $2A_\chi(2s)m_\chi^{-1}$ by reusing all the calculations of the appendix \ref{app A}, since $A_\chi(2s)\in I$ was the only hypothesis needed. We then deduce that $s^2+1$ is a square residue for at least one of an explicit finite set of modules.
This is also true for $D(b,s,k)$ (see \eqref{D McZim}) since it equals $s^2+1$ modulo a prime $q$ dividing $b$.
Therefore the ideal $(q)\subset\mathbb{Z}$ is not inert when extended in $\mathcal{O}_K$. 
If we assume that $h(D)=1$ (from now until the end of this section we will call $D(b,s,k)$ just $D$) there is in fact an algebraic integer of norm $q$ and, by the standard integral basis $(1,\sqrt{D})$, there are positive integers $A,B$ such that
$$|A^2-B^2D|=q.$$
Following the same lines of \cite[§3]{Biro2}, but extending the proof to the case of $D\not\equiv 1$ modulo $4$ (in \cite{Biro2} only the congruence $1$ modulo $4$ is treated), this will imply that
$$
\frac{Bq}{2B\sqrt{D}+\frac{q}{B\sqrt{D}}}\leq B^2\left|\frac{A}{B}-\sqrt{D}\right| \leq \frac{q}{\sqrt{D}}.
$$
and that $\frac{A}{B}$ is a convergent of $\sqrt{D}$. Calling $b_j$ the coefficients of the continued fraction, as explained by Schmidt \cite[p. 17]{Schmidt} it provides the following estimate
$$
\frac{1}{b_{j+1}+2}\leq B\left|A-B\sqrt{D}\right|\leq\frac{1}{b_{j+1}}.
$$
Combining the all these inequalities together we will force a contradiction. 

The last section is devoted to proving Proposition \ref{prop 1} and Proposition \ref{prop 2} following again the same strategy of \cite[§3]{Biro2}. When $D\equiv 1 $ modulo $4$ the continued fraction of $\frac{1+\sqrt{D}}{2}$ is derived from the one of $\sqrt{D}$ with little effort, see Lemma \ref{lemma divisione}.

\section{The family $n^2+1$}\label{family nn+1}

Taking inspiration from \cite{Biro1} and \cite{Biro3}, we look for a finite set of primes such that $n^2+1$ is a quadratic residue for at least one of them.
Just throughout this section we indicate by $D$ a squarefree number of the form $n^2+1$ with $n$ odd, so that
$\mathcal{O}_K=\mathbb{Z}[\sqrt{D}]$ and $d=4D$ is the fundamental discriminant.
It is not hard to verify that
$$\xi=n+\sqrt{n^2+1}$$
is the fundamental unit of $\mathcal{O}_K$.
Our first goal is to prove that

\begin{lem}\label{lemma fatto A}
If $D=n^2+1$, with $n$ an odd positive integer, is squarefree and $\chi$ is an odd primitive character of conductor $q$, where $q$ is a positive integer, then $\zeta_{P(K)}(s,\chi)$ can be extended meromorphically to the whole complex plane and it holds the formula
\begin{equation*}
\zeta_{P(K)}(0,\chi)=\frac{1}{q}A_\chi(2n);
\end{equation*}
see \eqref{def A} for the definition of $A_\chi$. Moreover if $(q,d)=1$ and $h(D)=1$ then
$m_\chi\neq0$ and $2A_\chi(2n)m_\chi^{-1}$ is an algebraic integer, where $m_\chi$ is defined in \eqref{def m}.
\end{lem}

\begin{rmk}
It is worth remarking that, differently from \cite{Biro1}, a factor 2 appears in front of $A_\chi$, but it does not cause any problem, indeed later we will work with some prime ideals $I$ over an odd integer prime $r$ containing $2A_\chi(2n)$, so we will have $A_\chi(2s)\in I$ which is a crucial condition in the proof of Theorem \ref{thm 2}.
\end{rmk}

To prove Lemma \ref{lemma fatto A} we need to extend the holomorphic function $\zeta_{P(K)}(s,\chi)$, originally defined for $\frak{Re}(s)>1$, to the whole complex plane and then to evaluate it at $s=0$. This will be done by writing the series defining $\zeta_{P(K)}(s,\chi)$ in a finite sum of series involving the Shintani zeta function, allowing us to apply \cite[Proposition (Shintani)]{Biro1}: this does not just give the meromorphic continuation, but also its value at $s=0$ expressed in term of Bernoulli polynomials.

\begin{defn}[Shintani zeta function]\label{Shintani Z}
Given a matrix $\begin{pmatrix}a&b\\c&d\end{pmatrix}$ with positive entries and $x>0,y\geq0$, the Shintani zeta function is defined as
\begin{equation}
\zeta\left(s,\begin{pmatrix}a&b\\c&d\end{pmatrix},x,y\right):=\sum_{n_1,n_2=0}^{\infty}(a(n_1+x)+b(n_2+y))^{-s}(c(n_1+x)+d(n_2+y))^{-s}.
\end{equation}
\end{defn}

\begin{proof}[Proof of Lemma \ref{lemma fatto A}]
We first want to describe in a suitable way the norm of a principal ideal by choosing a special generator.
Since $\xi\bar\xi=-1$, $\bar\xi<0$ and every principal ideal can be generated by a unique $\beta\in \mathcal{O}_K$ satisfying
\begin{equation}\label{ineq beta}
\beta>0,\quad \bar\beta>0,\quad 1\leq\beta/\bar\beta< \xi^4.
\end{equation}
As in \cite[Proof of Lemma 1]{Biro1}, $\beta$ can be written uniquely as $\beta=X+Y\xi^2$ with $X>0,Y\geq0$ in $\mathbb{Q}$ due to the fact that $(1,\xi^2)$ is a basis of $\mathbb{Q}(\sqrt{D})/\mathbb{Q}$ and to the inequalities in \eqref{ineq beta}; since $q$ is a positive integer, $X,Y$ can be in turn written uniquely as $X=qx+qn$ e $Y=qy+qm$ for non-negative integers $n,m$ such that $0<x\leq1,0\leq y<1$ (just perform the euclidean division).
At this point 
$$\beta\in \mathcal{O}_K \iff q(x+y\xi^2)\in \mathcal{O}_K.$$
Since $(1,\xi)$ is an integral basis, $C+D\xi$ with $0\leq C,D\leq q-1$ realizes a complete system of representatives for $\mathcal{O}_K/(q)$ and there exist unique $C,D$ such that
$$
qx+qy\xi^2\in C+D\xi+q\mathcal{O}_K,
$$
from which
$$x+y\xi^2-\frac{C+D\xi}{q}\in \mathcal{O}_K.$$
If we define 
\begin{equation}\label{def Q}
Q(C,D):=D^2-C^2-2nCD,
\end{equation}
and $a$ is the ideal generated by $X+Y\xi^2$, one has that
$$N(a)\equiv (C+D\xi)(\overline{C+D\xi})\equiv -Q(C,D)\quad (q).$$
It follows that
\begin{align*}
\zeta_{P(K)}(s,\chi)=&\sum_{(\beta)\in \mathcal{O}_K}\frac{\chi(N(\beta))}{N(\beta)^s}\\
=&\sum_{0\leq C,D\leq q-1}\sum_{(x,y)\in R(C,D)}\sum_{n,m\in\mathbb{N}}\frac{\chi(-Q(C,D))}{(X+Y\xi^2)^s(X+Y{\bar\xi}^2)^s}\\
=&\frac{-1}{q^{2s}}\sum_{0\leq C,D\leq q-1}\chi(Q(C,D))\sum_{(x,y)\in R(C,D)}\zeta\left(s,\begin{pmatrix}1&\xi^2\\ 1& \xi^{-2}\end{pmatrix},x,y\right),
\end{align*}
where the function in the last summand is the Shintani zeta function (see Definition \ref{Shintani Z}) and 
$$
R(C,D):=\{(x,y)\in\mathbb{Q}^2|0<x\leq1,0\leq y<1,x+y\xi^2-\frac{C+D\xi}{q}\in \mathcal{O}_K\}.
$$ 
From this it follows that
\begin{equation}\label{eq somma pesata}
\zeta_{P(K)}(0,\chi)=-\sum_{0\leq C,D\leq q-1}\chi(Q(C,D))\Sigma_{C,D},
\end{equation}
where
$$
\Sigma_{C,D}=\sum_{(x,y)\in R(C,D)}\left(B_1(x)B_1(y)+\frac{1}{4}\left( B_2(x)(\xi^2+\xi^{-2})+B_2(y)(\xi^2+\xi^{-2}) \right)\right)
$$
with $B_1(z)=z-\frac{1}{2}$ and $B_2(z)=z^2-z+\frac{1}{6}$ the first two Bernoulli polynomials.
An easy computation shows that
\begin{align*}
&\Sigma_{C,D}\\
=&\sum_{(x,y)\in R(C,D)}\left(\frac{2n^2+1}{2} (x+y)^2-2n^2(xy)-(n^2+1)(x+y)+\frac{4n^2+5}{12}\right).
\end{align*}
To calculate this sum it is necessary to understand better how $x,y$ are made; we will also show (see conditions \eqref{cond somma finita Shintani}) that $|R(C,D)|<\infty$ for fixed $C,D$ allowing us to get the analytic continuation and the value at $s=0$ for $\zeta_{P(K)}(s,\chi)$ from  Shintani zeta function. This can be done by changing basis over $\mathbb{Q}$ from $(1,\xi^2)$ to $(1,\xi)$. 
For $a,b\in\mathbb{Q}$, since $\xi=n+\sqrt{n^2+1}$ and $\xi^2=2n\xi +1$, one can check that
$$
\frac{a+b\xi}{q}=\frac{2na-b}{2nq}+\frac{b}{2nq}\xi^2
$$
and therefore, if we impose
$$
(x,y)=\left( \frac{1}{q}\left(a-\frac{b}{2n}\right),\frac{b}{2nq} \right),
$$
the elements of $R(C,D)$ are defined by the conditions
\begin{equation}\label{cond somma finita Shintani}
0<a-\frac{b}{2n}\leq q,\quad 0\leq \frac{b}{2n} <q,\quad \frac{(a-C)+(b-D)\xi}{q}\in \mathcal{O}_K;
\end{equation}
as a consequence $|R(C,D)|<\infty$. This also implies $a-C= qi$ and $b-D= qj$ with $i,j\in\mathbb{Z}$ and 
$$0<C+qi-\frac{D+qj}{2n}\leq q,\quad 0\leq \frac{D+qj}{2n} <q,$$
from which
$$0\leq j \leq 2n-1, \quad \frac{D+qj}{2nq}-\frac{C}{q}<i\leq\frac{D+qj}{2nq}-\frac{C}{q}+1.$$
With our choices the parameter $i$ is uniquely determined.
Indeed, if we define $A:=\lceil\frac{2nC-D}{q}\rceil$, then
\begin{align*}
&\frac{-(2nC-D)+qj}{2nq}<i\leq\frac{-(2nC-D)+qj}{2nq}+1 \\
\Longrightarrow\ &i=\begin{cases}0 & 0\leq j < A \\ 1 & A\leq j < 2n\end{cases}\Longrightarrow\ a=\begin{cases}C & 0\leq j < A \\ C+q & A\leq j < 2n\end{cases}.
\end{align*}
Calling $a_j:=a$ and $b_j:=b$ we can rewrite $\Sigma_{C,D}$:
$$\Sigma_{C,D}=\sum_{j=0}^{2n-1}\left(\frac{2n^2+1}{2q^2}a_j^2-\frac{2n^2}{q^2}\left(a_j-\frac{b_j}{2n}\right)\frac{b_j}{2n}-\frac{n^2+1}{q}a_j+\frac{4n^2+5}{12}\right).$$
At this point, through a long calculation,
\begin{align}
\begin{split}\label{calcolo sigma}
\Sigma_{C,D}=&A\left(1-\frac{C}{q}\right)+n\frac{(qA-2Cn+D)^2+D^2+2C^2}{2q^2}\\
-&\frac{(n+1)(qA-2nC+D)+(n-1)D+2nC}{2q}.
\end{split}
\end{align}
Since the first term of the right-hand side of \eqref{calcolo sigma} is the desired one, we just need to show that the last two summands, when we compute the weighted sum against the character in \eqref{eq somma pesata}, produce zero. As in \cite{Biro1} we introduce the transformation 
\begin{equation}\label{T}
T(C,D)=\left(D-2nC-q\left\lfloor \frac{D-2nC}{q} \right\rfloor,C\right).
\end{equation}
We will denote also as $(\widehat{C},\widehat{D})$.
This is a permutation of the elements of the set of couples $(C,D)$ with $0\leq C,D<q$. 
Indeed $0\leq D-2nC-q\lfloor\frac{D-2nC}{q}\rfloor<q$ is an integer and the transformation is invertible.
The following relations are straightforward:
\begin{equation*}
\widehat{C}-D+2nC=qA,\quad \widehat{C}=\widehat{\widehat{D}}.
\end{equation*}
Then
\begin{align*}
\Sigma_{C,D}=&A\left(1-\frac{C}{q}\right)+n\frac{\widehat{\widehat{D}}^2+\widehat{D}^2+\widehat{D}^2+D^2}{2q^2}\\
-&\frac{(n+1)(\widehat{\widehat{D}}+\widehat{D})+(n-1)(\widehat{D}+D)}{2q}
\end{align*}
and
$$
Q(C,D)\equiv -Q(\widehat{C},\widehat{D})\pmod{q},
$$
see \eqref{def Q} for the definition of the quadratic form $Q$.
Since $\chi$ is odd and when we apply $T$ the quadratic form $Q$ changes sign modulo $q$, one has $\chi(Q(C,D))=-\chi(Q(\widehat{C},\widehat{D}))$. Therefore the orbits of $T$ over which $\chi(Q(\cdot,\cdot))$ is not $0$ have even cardinality. This proves the first part of Lemma \ref{lemma fatto A}.

Now we prove the second one. Let us recall some basic facts about zeta functions and $L$-functions.
It is well known that $\zeta_K(s)=\zeta(s)L(s,\chi_d)$
and, with the same reasoning, one can prove that

\begin{equation}\label{eq ZL}
\zeta_K(s,\chi)=L(s,\chi)L(s,\chi\chi_d).
\end{equation}
Another classical fact we need (see \cite[chapter 9]{Davenport}) is that, for an odd primitive character $\chi$ of conductor $q$ one has

\begin{equation}\label{eq mL}
L(0,\chi)=-\frac{1}{q}\sum_{a=1}^q a\chi(a)\neq 0.
\end{equation}
If $(d,q)=1$ then $\chi\chi_d$ is primitive of conductor $qd$ and $\chi_d(-1)=1$ because $d>0$. Thus $m_\chi\neq 0$ by \eqref{eq mL} and 

\begin{equation}\label{eq ZL mL}
\zeta_K(0,\chi)=\frac{1}{q^2d}\cdot\sum_{a=1}^q a\chi(a)\cdot\sum_{b=1}^{qd} b\chi(b)\chi_d(b) =\frac{m_\chi}{q^2d}\cdot\sum_{b=1}^{qd} b\chi(b)\chi_d(b)
\end{equation}
by \eqref{eq ZL} and \eqref{eq mL}.

If we assume the hypotheses of the last statement of Lemma \ref{lemma fatto A}, precisely $h(D)=1$ and $(q,d)=1$, by \eqref{eq ZL mL} we get
$$
2A_\chi(2n)m_\chi^{-1}=\frac{2}{qd}\sum_{b=1}^{qd}b\chi(b)\chi_d(b)=\sum_{l=1}^q \chi(l)\sum_{x=0}^{d-1}\frac{2(l+xq)\chi_d(l+xq)}{qd}.
$$
To conclude it is sufficient to show that the last sum on $x$ is an integer for every $l$.
If we fix $l$ and vary $x$, then $l+xq$ spans all the classes modulo $d$ and $\chi_d$ is non trivial, thus
$$
\sum_{x=0}^{d-1}(l+xq)\chi_d(l+xq)\equiv l\sum_{x=0}^{d-1}\chi_d(l+xq)\equiv 0 \pmod{q}.
$$
For the divisibility by $d$, if we change variable setting $y=l+xq$, we see that
$$
2\sum_{x=0}^{d-1}(l+xq)\chi_d(l+xq)=2\sum_{y=0}^{d-1}y\chi_d(y)=\sum_{y=0}^d y\chi_d(y)+\sum_{y=0}^d (d-y)\chi_d(y)
$$
because $\chi_d(-1)=1$.
Thus
$$
2\sum_{x=0}^{d-1}(l+xq)\chi_d(l+xq)\equiv  0 \pmod{d}
$$
This complete the proof of Lemma \ref{lemma fatto A}.
\end{proof}

\begin{lem}\label{lemma 2.2}
If $0\neq\beta\in \mathcal{O}_K$ and $|\beta\bar\beta|<2n$ then $\beta$ is associated to an integer.
\end{lem}

\begin{proof}Let us write $\beta=e\xi^{-1}-f$ with $e,f\in\mathbb{Z}$. We can think $\xi^{-1}\leq |\beta|\leq 1$ and $e>0$.
$$|\bar\beta|=\left| -e\xi-f \right|=\left| e\xi+f \right|=\left| e(\xi+\xi^{-1})-\beta \right|\geq e(\xi+\xi^{-1})-1.$$
This implies
$$e- \xi^{-1} < e-\xi^{-1}+e\xi^{-2}\leq \xi^{-1}|\bar\beta| \leq|\beta\bar\beta|<2n,$$
where we also used $\xi^{-1}\leq |\beta|\leq 1$ and $\beta\bar\beta\in\mathbb{Z}$, providing a condition over $e$: $1\leq e \leq 2n-1.$
From
$e\xi^{-1}=e\cdot (-n+\sqrt{n^2+1})$
we deduce that $0<e\xi^{-1}<1$
and, thanks to the inequality $|\beta|\leq 1$, we deduce $d\in\{0,1\}$. If $d=1$ we have
$$
2n>|\beta\bar\beta|=|(e\xi^{-1}-1)(-e\xi-1)|=-e^2+2ne+1\geq 2n,
$$
which is a contradiction. Hence $d=0$ and $\beta$ is associated to an integer, as claimed.
\end{proof}

\begin{prop}\label{prop 2.3}
If $D=n^2+1$ with $n>1$ odd is squarefree and $h(D)=1$ then $D=2p$ where $p$ is an odd prime; if $r$ is an odd prime such that $r<2n$ then
$$\left(\frac{D}{r}\right)=-1.$$
\end{prop}

\begin{proof}
Let $p$ be the smallest odd prime divisor of $D$ and let us write $D=2pa$ with $a\geq p$. Then $2p^2\leq n^2+1$ and so $2<p<n.$
Because of $p|D$, in the ring of integers $(p)=(\beta)^2$ for $\beta\in \mathcal{O}_K$.
Since $|\beta\bar\beta|=p<2n$, by Lemma \ref{lemma 2.2} we deduce that $p$ is a square. This is absurd, so $D=2p$ with $p$ prime. Let $r$ be as in the statement; then $\left(\frac{D}{r}\right)\neq -1$ means that $(r)$ factors into two prime ideals and thus $|\beta\bar\beta|=r<2n,$
then $r$ is a square by Lemma \ref{lemma 2.2}.
It is absurd.
\end{proof}

As in \cite{Biro1}, if $m$ is a positive odd integer we call $U_m$ the set of the integers $a$ such that 
$$\left(\frac{a^2+1}{r}\right)=-1$$
for every prime divisor $r$ of $m$.
Let $\mathbb{Q}(\chi)$ be the field generated over $\mathbb{Q}$ by the values $\chi(a)$ and let $I$ be a prime ideal of $\mathbb{Q}(\chi)$ containing $m_\chi$, the algebraic integer defined in \eqref{def m}. 
Moreover if we write
$n=Pq+n_0$ with $0\leq n_0<q$ and make the substitution inside $A_\chi$
it follows that
\begin{align*}
A_\chi(2n)=A_\chi(2n_0)+2PB_\chi(2n_0),
\end{align*}
see \eqref{def B} for the definition of $B_\chi(n)$.
From \cite{Biro1} we recall the following condition:
the integer $q$ is odd, $r$ is an odd prime and there is an odd primitive character $\chi$ of conductor $q$ and a prime ideal $I$ of $\mathbb{Q}(\chi)$ lying above $r$ such that $m_\chi\in I$, but $I$ does not divide the ideal generated by $B_\chi(a)$ in the ring of integers of $\mathbb{Q}(\chi)$, if $a$ is any rational integer with $a\in U_q$.
\newline
We will be referring to it as ``condition $(*)$" or simply ``$(*)$".
If $(*)$ holds and $n_0\in U_q$, then by Lemma \ref{lemma fatto A}
$$
P\equiv - \frac{A_\chi(2n_0)}{2B_\chi(2n_0)}\pmod{I}
$$
and, consequently,
$$
n\equiv n_0- q\frac{A_\chi(2n_0)}{2B_\chi(2n_0)}\pmod{I}.
$$
As explained just before Remark \ref{rmk 2.1} the choices of $q$, $n_0$ and of the character $\chi$ and the validity of the condition $(*)$ will force $n$ to stay in a precise residue class modulo $r$. 
As we will show in appendix \ref{app A}, the choices of characters and prime ideals made by Biró in \cite{Biro1} for Yokoi's family still work for the family $n^2+1$.
Thus for each of the four characters of \cite[§4]{Biro1} we calculate the values of $A_\chi(2n_0)$ and $B_\chi(2n_0)$ modulo $I$ for every residue class $n_0$ modulo $q$. We then compute the residue class of $n_0$ modulo $r$ and delete those classes such that $n_0^2+1$ is a square modulo $r$ (see condition $(*)$). Repeating this process we are able to prove Theorem \ref{thm 2}.
In appendix \ref{app B}, for the sake of completeness, we exhibit an example of the code used for the calculations.

As a conclusion of this section, we add that Theorem \ref{thm 2} and Proposition \ref{prop 2.3} imply Corollary \ref{cor 1 thm 2} and Corollary \ref{cor 2 thm 2}.

\section{Proof of Theorem \ref{thm 1}} \label{proof thm 1}

Throughout this section we assume that $q|b$, $D$ is as in \eqref{D McZim} squarefree with $b+s$ odd. The latter condition is equivalent to $D\equiv 2$ modulo $4$. 
If we call $\tau:=(4bs+1)$, by \cite[Theorem 6]{McZimmer} we get
\begin{equation}\label{expansion 1}
\sqrt{D}=[b\tau^k+s,\overline{2b,2b\tau^{k-1},2b\tau,2b\tau^{k-2}, \ldots, 2b\tau^{k-1},2b,2b\tau^k+2s}].
\end{equation} 
In order to have information about $\zeta_{P(K)}$ we apply \cite[Theorem 1]{BiroGranville}.
If $\sqrt{D}=[a_0,\overline{a_1,\ldots,a_l}]$
let us set $\omega=\sqrt{D}$, $\alpha=\omega-a_0$, $\frac{p_j}{q_j}=[0,a_1,\ldots,a_j]$, $\alpha_j=p_j-q_j\alpha$ and $\alpha_0=-\alpha$.
By \eqref{expansion 1} it follows that 
$$
p_1=1,\quad q_1=2b,\quad p_2=2b\tau^{k-1},\quad q_2=4b^2\tau^{k-1}+1
$$
and the coefficients $p_j$ and $q_j$ are alternatively equivalent to $0$ or $1$ modulo $q$. Indeed we have
\begin{align*}
&p_1\equiv p_3\equiv\ldots\equiv p_{2k-1}\equiv q_2\equiv q_4\equiv\ldots\equiv q_{2k}\equiv 1\pmod{q},\\
&p_2\equiv p_4\equiv\ldots\equiv p_{2k}\equiv  q_1\equiv q_3\equiv\ldots\equiv q_{2k-1}\equiv 0\pmod{q},
\end{align*}
with exception of the last convergent for which we have
$$
p_{2k+1}\equiv 1\pmod{q},\quad q_{2k+1}\equiv 2s\pmod{q}.
$$

As in \cite{Biro2} we introduce the quadratic forms
$$Q_j(x,y)=(\alpha_{j-1}x+\alpha_jy)(\bar\alpha_{j-1}x+\bar\alpha_jy)$$
and, if we write
$$Q_j(x,y)=A_jx^2+B_jxy+C_jy^2,$$
we get
\begin{align*}
& A_1=\alpha\bar\alpha,\qquad B_1=2q_1\alpha\bar\alpha-p_1(\alpha+\bar\alpha),\\
&A_j=p_{j-1}^2-p_{j-1}q_{j-1}(\alpha+\bar\alpha)+q_{j-1}^2\alpha\bar\alpha \qquad(j>1),\\
&B_j=2p_{j-1}p_j+2q_{j-1}q_j\alpha\bar\alpha-(p_{j-1}q_j+p_jq_{j-1})(\alpha+\bar\alpha)\qquad (j>1),\\
&C_j=p_j^2-p_jq_j(\alpha+\bar\alpha)+q_j^2\alpha\bar\alpha\qquad (j\geq 1).
\end{align*}
From the congruences $-(\alpha+\bar\alpha)\equiv 2s, \alpha\bar\alpha\equiv -1$ modulo $q$,
we derive the next congruences modulo $q$
$$A_j\equiv(-1)^j,\ B_j\equiv2s\ (j\leq 2k), B_{2k+1}\equiv-2s,\ C_j\equiv(-1)^{j-1}.$$

\begin{prop}\label{prop 4.1}
Call by $d$ the discriminant of $\mathbb{Q}(\sqrt{D})$. If $(q,d)=1$ and $\chi$ is an odd character of conductor $q$, then $\zeta_{P(K)}$ extends meromorphically to the whole complex plane and
\begin{align*}
\zeta_{P(K)}(0,\chi)=&\frac{2}{q^2}\sum_{1\leq u,v\leq q-1}uv\chi(u^2+2suv-v^2)+\\
+&\left(\frac{d}{q}\right)\frac{\chi(-d)\tau(\chi)^2L(2,\overline\chi^2)}{\pi^2}\left(2b\tau^k+2s+\frac{\tau^k-1}{s}\right).
\end{align*}
\end{prop}

\begin{proof}
The thesis follows from \cite[Theorem 1]{BiroGranville} indeed, if $A_j,B_j,C_j$ are as above,
\begin{align*}
\zeta_{P(K)}(0,\chi)=&\frac{2}{q^2}\sum_{j=1}^l\sum_{1\leq u,v\leq q-1}uv\chi((-1)^j(A_ju^2+B_juv+C_jv^2))\\
+&\left(\frac{d}{q}\right)\frac{\chi(-d)\tau(\chi)^2L(2,\overline\chi^2)}{\pi^2}\sum_{j=1}^l a_j\overline\chi((-1)^jA_j).
\end{align*}
Now using the congruences obtained just before Proposition \ref{prop 4.1} we get
\begin{align*}
&\zeta_{P(K)}(0,\chi)\\
=&\frac{2}{q^2}\sum_{j=1}^{2k}\sum_{1\leq u,v\leq q-1}uv\chi(u^2+(-1)^j2suv-v^2)\\
+&\frac{2}{q^2}\sum_{1\leq u,v\leq q-1}uv\chi(u^2+2suv-v^2)+\left(\frac{d}{q}\right)\frac{\chi(-d)\tau(\chi)^2L(2,\overline\chi^2)}{\pi^2}\sum_{j=1}^l a_j\\
=&\frac{2}{q^2}\sum_{1\leq u,v\leq q-1}uv\chi(u^2+2suv-v^2)\\
+&\left(\frac{d}{q}\right)\frac{\chi(-d)\tau(\chi)^2L(2,\overline\chi^2)}{\pi^2}\left(2b\tau^k+2s+\frac{\tau^k-1}{s}\right).
\end{align*}
\end{proof}

A reasoning similar to that of Lemma 2.2. of \cite{Biro2} gives the
\begin{lem}\label{lemma 4.2.}
Under the same hypotheses of the previous proposition
$$\chi(-d)\left(\frac{d}{q}\right)\frac{\tau(\chi)^2L(2,\overline\chi^2)}{\pi^2}=\sum_{u,v}\left(\frac{v^2}{q^2}-\frac{v}{q}\right)\chi(u^2+2suv-v^2).$$
\end{lem}

\begin{lem}\label{lemma 4.3.}
Let $\chi$ be an odd primitive character of conductor $q$ with $(q,d)=1$ and of order greater than two. If $A_\chi(n)$ is defined as in \eqref{def A}, then
\begin{align}\label{eq lemma 5.3.}
\begin{split}
\frac{1}{q}A_\chi(2s)=&\frac{2}{q^2}\sum_{1\leq u,v \leq q-1}uv\chi(u^2+2suv-v^2)\\
+&2s\sum_{0\leq u,v \leq q-1}\left(\frac{v^2}{q^2}-\frac{v}{q}\right)\chi(u^2+2suv-v^2).
\end{split}
\end{align}
\end{lem}

\begin{proof}
Recall the transformation $(\widehat{C},\widehat{D})$ defined in \eqref{T}; one can easily verify that
$$
A\frac{(C-q)}{q}=\frac{1}{q^2}\left(  (C-q)(2sC-D)+\widehat{C}\widehat{D}-q\widehat{C}\right),
$$
where $A:=\lceil\frac{2nC-D}{q}\rceil$. Moreover, as explained in the end of the proof of Lemma \ref{lemma fatto A}, the orbits of $T$ on which $\chi$ is not trivial have even cardinality since $\chi(C^2+2sCD-D^2)$ changes sign with every application of $T$. Hence
\begin{align*}
&\frac{A_\chi(2s)}{q}=\sum_{0\leq C,D\leq q-1}\chi(D^2-C^2-2sCD)\frac{A(C-q)}{q}\\
=&\sum_{0\leq C,D\leq q-1}\chi(D^2-C^2-2sCD)\frac{1}{q^2}\left(  (C-q)(2sC-D)+\widehat{C}\widehat{D}-q\widehat{C}\right)
\end{align*}
and, since 
$$
\sum_{0\leq C,D\leq q-1}\chi(D^2-C^2-2sCD)\left( \widehat{C}\widehat{D}-q\widehat{C}\right)
$$
is equal to
$$
\sum_{0\leq C,D\leq q-1}\chi(D^2-C^2-2sCD)\left( -CD+qC \right),
$$
then
$$
\frac{A_\chi(2s)}{q}=\sum_{0\leq C,D\leq q-1}\chi(D^2-C^2-2sCD)\frac{1}{q^2}\left(  (C-q)(2sC-D)-CD+qC \right).
$$
We now show that the difference between
$\frac{1}{q}A_\chi(2s)$ and \eqref{eq lemma 5.3.} is zero.
Writing $(u,v)=(q-C,D)$ in \eqref{eq lemma 5.3.} one gets
\begin{align*}
&\frac{2}{q^2}\sum_{1\leq C,D\leq q-1}(q-C)D\chi(D^2-2sCD-C^2)\\
+&2s\sum_{0\leq C,D \leq q-1, 1\leq C}\left(\frac{(q-C)^2}{q^2}-\frac{q-C}{q}\right)\chi(D^2-2sCD-C^2)\\
-&\frac{1}{q^2}\sum_{0\leq C,D\leq q-1}\chi(D^2-C^2-2sCD)\left(  (C-q)(2sC-D)-CD+qC \right)
\end{align*}
and, after other calculations, the above becomes
\begin{align}
\begin{split}\label{eq 2 lemma 5.3.}
&\frac{1}{q}\sum_{1\leq C,D\leq q-1}D\chi(D^2-2sCD-C^2)-\frac{1}{q}\sum_{1\leq D\leq q-1}D\chi(D^2)\\
-&\frac{1}{q}\sum_{0\leq C,D\leq q-1}C\chi(D^2-2sCD-C^2).
\end{split}
\end{align}
Changing variables $(C,D)\mapsto (-C,q-D)$ in the first sum of \eqref{eq 2 lemma 5.3.}, $D\mapsto q-D$ in the second one and $(C,D)\mapsto(q-C,-D)$ in the third one, we see that \eqref{eq 2 lemma 5.3.} is zero.
\end{proof}

\begin{lem}\label{lemma 4.4}
Let $D$ be squarefree with $b,s$ verifying the hypotheses of Theorem \ref{thm 1}. Let $\chi$ be as above with conductor $q$ ($(q,d)=1$) and of order greater than two. Let us assume that $q|b$.
Then $$q\zeta_{P(K)}(0,\chi)-A_\chi(2s)$$
is equal to $\frac{b}{q}$ times an algebraic integer.
\end{lem}

\begin{proof}
By what proved in Proposition \ref{prop 4.1}, Lemma \ref{lemma 4.2.} and Lemma \ref{lemma 4.3.}
\begin{align*}
&q\zeta_{P(K)}(0,\chi)-A_\chi(2s)\\
=&\frac{b}{q}\sum_{0\leq u,v\leq q-1}(v^2-qv)\chi(u^2+2suv-v^2)\left(2\tau^k+\frac{\tau^k-1}{bs}\right)
\end{align*}
and $bs$ divides $\tau^k-1$ is an integer.
\end{proof}

Equations \eqref{eq mL} and \eqref{eq ZL mL}, by reasoning as in the last part of the proof of Lemma \ref{lemma fatto A}, give the following
\begin{lem}\label{lem 2.4.}
Let $D$ be squarefree with $b,s$ as above. Let $\chi$ be an odd primitive character of conductor $q$ with $(q,d)=1$. Let us assume that $h(D)=1$.
Then $m_\chi\neq 0$ and $$2q\zeta_K(0,\chi)m_\chi^{-1}$$
is an algebraic integer.
\end{lem}

\begin{lem}
If the hypotheses of Lemma \ref{lem 2.4.} hold with $\chi$ being of order greater than two and if moreover there exists a prime ideal $I$ of $\mathbb{Q}(\chi)$ containing $m_\chi$ and an odd prime $r$ such that $r|\frac{b}{q}$, then $A_\chi(2s)\in I$.
\end{lem}

\begin{proof}
$\zeta_K(0,\chi)\in \mathbb{Q}(\chi)$
and $m_\chi \in I$, by Lemma \ref{lem 2.4.}, imply $2q\zeta(0,\chi)\in I.$
Since by Lemma \ref{lemma 4.4}
$$2q\zeta(0,\chi)-2A_\chi(2s)\in\frac{b}{q}\mathcal{O}_K$$
and $r|\frac{b}{q}$, we have $2q\zeta(0,\chi)-2A_\chi(2s)\in I$.
It follows that $2A_\chi(2s)\in I$
and, as $I$ is prime and it is not above $2$, $A_\chi(2s)\in I$. 
\end{proof}

In the previous section we exhibited some characters $\chi$ and prime ideals $I$. If we repeat the reasoning of section \ref{family nn+1}, we can show that if $A_\chi(2s)\in I$ for all these characters and ideals, then
$s^2+1$ is a square for at least one of the moduli $q=5,7,41,61,1861$.
We point out that here it is not necessary for $s$ to be odd.
We proved the

\begin{prop}\label{teorema num class}
Let $b,s,k>0$ such that $b+s$ is odd and $D$ is squarefree, let us assume $h(D)=1$. Suppose moreover that $41\cdot61\cdot175\cdot1861|b$. Then $D$ is a square for at least one of the following moduli $q=5,7,41,61,1861.$
\end{prop}

We now prove Theorem \ref{thm 1}.
\begin{proof}[Proof of Theorem \ref{thm 1}]
Let $b$ be positive and $q\in\{5,7,41,61,1861\}$ such that $D$ is a square modulo $q$. Then the ideal $(q)$ factors into two prime ideals, not necessarily distinct. That, together with the hypothesis $h(D)=1$, implies the existence of $\omega\in\mathcal{O}_K$ of norm $q$.
Since $D\not\equiv1$ modulo $4$ let $A,B\in\mathbb{Z}$ be such that 
$$\omega=A-B\sqrt{D}.$$
The equality $|\omega\bar\omega|=q$ is equivalent to
$$|A^2-B^2D|=q,$$
that in turn implies $B\neq0$ and $(A,B)=1$. We can assume $A,B>0$.
One can verify that $|\omega|\leq\frac{q}{B\sqrt{D}}$ and therefore 
$B|\omega|<\frac{1}{2}$.
This implies that $\frac{A}{B}$ is a convergent of $\sqrt{D}$. Moreover 
by the triangle inequality applied to $|\bar\omega|=|\omega+2B\sqrt{D}|$, it follows that
$$
|\bar\omega|\leq 2B\sqrt{D}+\frac{q}{B\sqrt{D}}
$$
and in turn
\begin{align}\label{diseq1}
\frac{Bq}{2B\sqrt{D}+\frac{q}{B\sqrt{D}}}\leq B|\omega| \leq \frac{q}{\sqrt{D}}.
\end{align}
Instead if we write $\frac{A}{B}=[b_0,b_1,\ldots,b_j]$ ($\frac{A}{B}$ is a convergent) for some $j$ with notation as in Lemma \ref{lemma divisione}, by the third displayed formula in  \cite[p, 17]{Schmidt} we get
\begin{align}\label{diseq2}
\frac{1}{b_{j+1}+2}\leq B|\omega|\leq\frac{1}{b_{j+1}}.
\end{align}
The inequalities \eqref{diseq1} e \eqref{diseq2} must hold at the same time.
For the inequality \eqref{diseq2}, due to the periodicity of the continued fraction, we can think $0\leq j\leq l-1$ where $l$ is the period of the continued fraction of $\sqrt{D}$.

\begin{itemize}
   \item[a)] Let $j$ be such that $0\leq j\leq 2k-1$.
Thus $1\leq b_{j+1}\leq 2b(4bs+1)^{k-1}$
and, using the upper bound from \eqref{diseq1} and the lower bound from \eqref{diseq2}, we conclude that 
$$
\frac{1}{2b(4bs+1)^{k-1}+2}\leq \frac{q}{\sqrt{D}}.
$$ 
Then
\begin{equation*}
b(4bs+1)^k+s=\lfloor\sqrt{D}\rfloor<\sqrt{D}\leq 2q[b(4bs+1)^{k-1}+1].
\end{equation*}

   \item[b)]
If $j=2k$ using the lower bound from \eqref{diseq1} and the upper bound from \eqref{diseq2} we get 
$$
\frac{Bq}{2B\sqrt{D}+\frac{q}{B\sqrt{D}}}\leq \frac{1}{b_{2k+1}}.
$$
This implies
$$
(q-1)\frac{2b(4bs+1)^k+2s}{q}\leq \frac{2}{q}+\frac{1}{B^2\sqrt{D}}<2.
$$
\end{itemize}
We get then
$$
b< \frac{b(4bs+1)^k+s}{2b(4bs+1)^{k-1}+2}\leq q
$$
or
$$
(q-1)\frac{2b(4bs+1)^k+2s}{q} <2,$$
but they are both a contradiction because relatively to the former $q|b$, while for the latter $q\geq2$ and $k\geq1$.

If $h(D)=1$, by Proposition \ref{teorema num class} and the previous discussion we deduce that $b=0$, hence $D=s^2+1$ with $s$ odd. We complete the proof by Corollary \ref{cor 2 thm 2}.
\end{proof}

\section{Proof of Propositions \ref{prop 1} and \ref{prop 2}} \label{proof prop 1}
Let us assume $D$ is as in \eqref{D McZim} so that the continued fraction of $\sqrt{D}$ is as in \eqref{expansion 1}. We recall that we defined $\tau:=4bs+1$.
As we want to deal with both the cases $D\equiv 1$ modulo $4$ and $D\not\equiv 1$ modulo $4$, it is useful to derive the continued fraction of $\frac{1+\sqrt{D}}{2}$ from that of $\sqrt{D}$. 

\begin{lem}\label{lemma divisione}
Let $x$ be a real number with continued fraction expansion
$$x=[a_0,\overline{a_1,a_2,\ldots,a_l}].$$
Assume that $a_i\geq3$ for $i\geq 1$ and $a_i$ even for $i\geq0$, then

\begin{equation*}
\frac{1+x}{2}=\left[\frac{a_0}{2},\overline{1,1,\frac{a_1-2}{2},1,1,\frac{a_2-2}{2},1,1,\ldots,\frac{a_l-2}{2}}\right].
\end{equation*}
\end{lem}

\begin{proof}
We just sketch the proof. Since $a_0$ is even, it follows that
$\cfrac{1+x}{2}=\cfrac{a_0}{2}+\cfrac{1}{\cfrac{2\alpha_1}{\alpha_1+1}},$
where $\alpha_j:=[a_j,a_{j+1},\ldots]$.
Since $\alpha_1>3$, we get
$$\frac{1+x}{2}=\frac{a_0}{2}+\cfrac{1}{1+\cfrac{1}{1+\cfrac{1}{\cfrac{\alpha_1-1}{2}}}}.$$
Therefore
$$
\frac{1+x}{2}=\left[\frac{a_0}{2},1,1,\frac{\alpha_1-1}{2}\right]
$$
and as $\lfloor \alpha_1-1 \rfloor$ is odd and $\alpha_1>3$ one can go on in the same way proving the lemma.
\end{proof}
In particular if $D$ is as in \eqref{D McZim} ans $D\equiv 1$ modulo $4$ with $b\geq2$ and $k\geq1$, then
$$
\frac{1+\sqrt{D}}{2}=[b_0,\overline{b_1,b_2,\ldots,b_{3(2k+1)}}]
$$
with

\begin{align}\label{esp lemma divisione}
\begin{split}
&b_{3(2i-1)}=b\tau^{i-1}-1,\quad b_{3(2i)}=b\tau^{k-i}-1 \qquad (1\leq i\leq k)\\
&b_0=\frac{b\tau^k+s}{2},\quad b_{3(2k+1)}=2b_0-1,\quad \mbox{and}\quad b_i=1\  \mbox{otherwise.}
\end{split}
\end{align}

Now let $(b,s)>2$ and $q$ be a prime dividing $(b,s)$.
Since $D\equiv 1$ modulo $q$ it is then a quadratic residue modulo $q$ and thus the ideal $(q)$ factors in the product of $2$ ideals in $\mathcal{O}_K$ if $q$ is odd. Actually $(b,s)=2^t$ with $t\geq2$ implies that also $(2)$ is not inert. Therefore we can always choose $q$ so that $(q)$ factors into two prime ideals of norm $q$ and, if we add the hypothesis $h(D)=1$, then there exists $\omega\in \mathcal{O}_K$ such that $|\omega\bar\omega|=q$; we want to use the property of the continued fractions of $\sqrt{D}$ or $\frac{1+\sqrt{D}}{2}$ to get a contradiction.  

\begin{proof}[Proof of Proposition \ref{prop 1}]
In the following we assume $h(D)=1$ and $(b,s)>2$ and we distinguish two cases based on if $D\equiv1$ modulo $4$ or not, reasoning respectively as \cite[§3]{Biro2} and as in the proof of Theorem \ref{thm 1}. In each case we are led to two possible inequalities for both cases:
\begin{equation}\label{eq 1 def prop 1}
b\tau^k+s<\sqrt{D}\leq 2q[b\tau^{k-1}+1].
\end{equation}
and
\begin{equation}\label{eq 2 def prop 1}
(q-1)\frac{b_l}{q} <2,
\end{equation}
where $b_l$ is the last term of the period of the continued fraction we are dealing with.
By \eqref{eq 1 def prop 1} and $q|(b,s)$ it follows that
$$
2q[b\tau^{k-1}+1]\leq 2bs\tau^{k-1}+2bs\leq 4bs\tau^{k-1}<\tau^k
$$ 
and thus $b\tau^k+s<\tau^k$ which is absurd. 
Inequality \eqref{eq 2 def prop 1} also gives a contradiction because $q\geq2$ and $k\geq1$.
This proves Proposition \ref{prop 1}.
\end{proof}

\begin{proof}[Proof of Proposition \ref{prop 2}]
Let us suppose $h(D)=1$ and $q$ is the smallest prime dividing $\tau$. Since $\tau$ is odd and not a prime, $q$ is odd and $q\leq \tau^{1/2}$. In order to settle Proposition \ref{prop 2}, we need to find out the continued fraction of $\frac{1+\sqrt{D}}{2}$ also when $b=1$ because equations \ref{esp lemma divisione} assume $b\geq 2$. We omit the proof since it goes as the one of Lemma \ref{lemma divisione}.

\begin{lem}\label{lemma divisione 2}
Being $D$ as in \eqref{D McZim}, if $b=1$, $k\geq2$ and $D\equiv 1$ modulo $4$ then
$$
\frac{1+\sqrt{D}}{2}=[b_0,\overline{b_1,b_2,\ldots,b_{6k-1}}]
$$
with
\begin{align*}
&b_{6(i-1)+7}=\tau^{i}-1,\quad b_{6(i-1)+4}=\tau^{k-i}-1 \qquad (1\leq i\leq k-1)\\
&b_0=\frac{\tau^k+s}{2},\quad b_{6k-1}=2b_0-1,\quad b_2,b_{6k-3}=2,\quad \mbox{and}\quad b_i=1\  \mbox{otherwise.}
\end{align*}
If instead $k=1$ then
$$
\frac{1+\sqrt{D}}{2}=[\frac{5s+1}{2},\overline{1,2,2,1,5s}].
$$
\end{lem}
If $q$ is a prime as above $D$ is obviously a square modulo $q$ and therefore we can proceed as in the proof of Proposition \ref{prop 1} getting, for each possible case ($D\equiv1$ or $D\not\equiv1$ modulo $4$), the inequalities \eqref{eq 1 def prop 1} and
\begin{equation}\label{eq 2 def prop 2}
b\tau^k+s<5.
\end{equation}
Inequality \eqref{eq 2 def prop 2} is actually quite rough but in that form it is valid for every case. 
Just for $b=k=1$ and $D\equiv 1 $ modulo $4$ \eqref{eq 1 def prop 1} is replaced by $\tau+s<8q$. Clearly \eqref{eq 2 def prop 2} is always absurd while $\tau+s<8q$ is absurd if $\tau\geq 64$ since $q\leq\tau^{1/2}$, so let us divide the rest of the proof into cases.

\begin{itemize} 
   \item[1)] 
Assume $b,k=1$. If $\tau\geq 64$ we immediately get an absurd for every class of $D$ modulo $4$, so if $h(D)=1$ then
$$
\tau \in [9,21,25,33,45,49,57]
$$
or equivalently
$$
s\in [2,5,6,8,11,12,14]
$$
and we check by PARI/GP that all class numbers are greater than one.

   \item[2)] Assume not both $b$ and $k$ are $1$ so that we can focus just on \eqref{eq 1 def prop 1}. If we define $a:=\log_9{6}$
we can show that
\begin{equation}\label{eq 1 def prop 1 bis}
\tau^a\leq \frac{b\tau^k+s}{b\tau^{k-1}+1}
\end{equation}
for all $k$; indeed this is equivalent to $1\leq b\tau^{k-1}(\tau^{1-a}-1)+s\tau^{-a}$
and 
$$
b\tau^{k-1}(\tau^{1-a}-1)+s\tau^{-a}\geq 2(\tau^{1-a}-1)\geq 2(9^{1-a}-1)=1.
$$
Since $\tau\geq9$, on the other hand we have that $2q\leq 2\tau^{1/2}\leq \tau^a$
and so, by \eqref{eq 1 def prop 1} and \eqref{eq 1 def prop 1 bis}, $\tau^a<2q\leq \tau^a$.
\end{itemize}
The proof is complete.
\end{proof}


\addresshere

\clearpage

\appendix

\section{Outcome of computations}\label{app A}

The first character $\chi_1$ is of conductor $175$, $r$ is set equal to $61$ and $I=(61,i\omega\xi-10)$, where $\omega$ is a primitive third root of unity and $\xi$ a primitive fifth root of unity. It is straightforward to check that $m_{\chi_1}\in I$ (this is actually stated in \cite{Biro1} but we preferred to repeat all the calculations with PARI/GP).
Now we write down the residue classes $n_0$ modulo 175 which are in $U_{175}$ (for the others we know by definition of $U_{175}$ that ${n_0}^2+1$ is a quadratic residue modulo $5$ or $7$) together with
the corresponding values of $A_{\chi_1}$, $B_{\chi_1}$, the class of $n$ modulo $r$ ($r=61$) and, eventually, the Kronecker symbol 
$\left(\frac{{n_0}^2+1}{r}\right)$.
We get 

{\centering
\begin{longtable}{c|c|c|c|c}
\hline
$n_0$ & $A_{\chi_1}(2n_0)$ & $B_{\chi_1}(2n_0)$ & $n$ mod  $61$ & $\left(\frac{\cdot}{\cdot}\right)$\\
\hline
4& 0& 33& 4& -1\\
9& 34& 44& 1& -1\\
11& 34& 53& 55& -1\\
16& 1& 44& 5& -1\\
19& 40& 30& 4& -1\\
24& 20& 50& 50& 0\\
26& 14& 32& 43& 1\\
31& 42& 24& 38& 1\\
39& 27& 53& 56& -1\\
44& 11& 30& 17& 1\\
46& 1& 33& 11& 0\\
51& 26& 32& 39& 1\\
54& 59& 26& 49& -1\\
59&30& 23& 43& 1\\
61& 38& 51& 58& -1\\
66&38& 23& 1& -1\\
74& 26& 50& 59& 1\\
79& 26& 26& 22& 1\\
81& 24& 24& 24& -1\\
86& 51& 51& 29& 1\\
89& 51& 51& 32& 1\\
94& 24& 24& 37& -1\\
96& 26& 26& 39& 1\\
101& 13& 50& 2& 1\\
109& 8& 23& 60& -1\\
114&3& 51& 3& -1\\
116& 16& 23& 18& 1\\
121& 54& 26& 12& -1\\
124& 38& 32& 22& 1\\
129&4& 33& 50& 0\\
131& 49& 30& 44& 1\\
136&18& 53& 5& -1\\
144& 6& 24&23& 1\\
149& 50& 32& 18& 1\\
151& 19& 50& 11& 0\\
156& 20& 30& 57& -1\\
159& 26& 44& 56& -1\\
164& 11& 53& 6& -1\\
166& 54& 44& 60& -1\\
171& 5& 33& 57& -1\\
\end{longtable}\par
}

\medskip
We are left to consider the following residues modulo $175$:
\begin{verbatim}
[4, 9, 11, 16, 19, 39, 54, 61, 66, 81, 94, 109, 114, 121, 
136, 156, 159, 164, 166, 171]
\end{verbatim}

The second character $\chi_2$ is also of conductor $175$, $r$ is set to be equal to $1861$ and $I=(1861,i\omega\xi-173)$, where again $\omega$ is a primitive third root of unity and $\xi$ is a primitive fifth root of unity. As before we see that $m_{\chi_2}$ is in $I$.
This time we get
\medskip

{\centering
\begin{longtable}{c|c|c|c|c}
\hline
$n_0$ & $A_{\chi_2}(2n_0)$ & $B_{\chi_2}(2n_0)$ & $n$ mod  $1861$ & $\left(\frac{\cdot}{\cdot}\right)$\\
\hline
4& 0& 1121& 4& 1\\
9& 1254& 1060& 1073& 1\\
11& 60& 1588& 746& 1\\
16& 135& 1060& 1484& -1\\
19& 1633& 1397& 1091& 1\\
39& 344& 1588& 531& -1\\
54& 1808& 1720& 1044& -1\\
61& 114& 1389& 1332& -1\\
66& 1669& 1102& 88& 1\\
81& 1498& 1498& 924& -1\\
94& 1498& 1498& 937& -1\\
109& 535& 1102& 1773& 1\\
114& 803& 1389& 529& -1\\
121& 1632& 1720& 817& -1\\
136& 971& 1588& 1330& -1\\
156& 1161& 1397& 770& 1\\
159& 124& 1060& 377& -1\\
164& 1255& 1588& 1115& 1\\
166& 866& 1060& 788& 1\\
171& 381& 1121& 1857& 1\\
\end{longtable}\par
}
\medskip

We have to check the remaining classes:
\newline

{\centering
\begin{longtable}{c|c|c}
\hline
$n$ mod $175$ & $n$ mod $61$ & $n$ mod $1861$\\
\hline
16 & 5 & 1484\\
39 & 56 & 531\\
54 & 49 & 1044\\
61 & 58 & 1332\\
81 & 24 & 924\\
94 & 37 & 937\\
114 & 3 & 529\\
121 & 12 & 817\\
136 & 5 & 1330\\
159 & 56 & 377\\
\end{longtable}\par
}
\medskip

to study which we use the other two characters.
The third one, $\chi_3$, is of conductor $61$, $r$ is equal to $1861$ and $I=(1861,i\omega\xi-1833)$, where $\omega$ and $\xi$ are as above.
One checks that $m_{\chi_3}$ is in $I$.
We obtain
\newline

{\centering
\begin{longtable}{c|c|c|c|c}
\hline
$n_0$ & $A_{\chi_3}(2n_0)$ & $B_{\chi_3}(2n_0)$ & $n$ mod  $1861$ & $\left(\frac{\cdot}{\cdot}\right)$\\
\hline
5& 1150& 616& 1371& -1\\
56& 82& 616& 490& -1\\
49& 1153& 663& 1631& 1\\
58& 1043& 1000& 1555& 1\\
24& 1347& 1347& 924& -1\\
37& 1347& 1347& 937& -1\\
3& 957& 1000& 306& 1\\
12& 173& 663& 230& 1\\
5& 1150& 616& 1371& -1\\
56& 82& 616& 490& -1\\
\end{longtable}\par
}
\medskip

We have to study the following classes
\newline

{\centering
\begin{longtable}{c|c|c|c}
\hline
$n$ mod $175$ & $n$ mod $61$ & $n$ mod $1861$ & $n$ mod 1861\\
\hline
16 & 5 & 1484 & 1371\\
39 & 56 & 531 & 490\\
81 & 24 & 924&924\\
94 & 37 & 937&937\\
136 & 5 & 1330&1371\\
159 & 56 & 377&490\\
\end{longtable}\par
}
\medskip

and, by requiring the compatibility of the classes modulo $1861$, we are left with
\newline

{\centering
\begin{tabular}{c|c|c}
\hline
$n$ mod $175$ & $n$ mod $61$ & $n$ mod $1861$ \\
\hline
81 & 24 & 924\\
94 & 37 & 937\\
\end{tabular}\par
}
\medskip

for which we use the fourth character.
The last one, $\chi_4$, is of conductor $61$, $r$ is equal to $41$ and $I=(41,i\xi-33)$, where $\xi$ is has above. We see again that $m_{\chi_4}$ is in $I$.
Through PARI/GP one gets
\newline

{\centering
\begin{tabular}{c|c|c|c|c}
\hline
$n_0$ & $A_{\chi_4}(2n_0)$ & $B_{\chi_4}(2n_0)$ & $n$ mod  $41$ & $\left(\frac{\cdot}{\cdot}\right)$\\
\hline

24& 13& 13& 14& 1\\
37& 13& 13& 27& 1\\

\end{tabular}\par
}
completing the proof of Theorem \ref{thm 2}.

\section{Example of code}\label{app B}

With the aim of giving an example of the code used, we just write the one relative to the first character $\chi_1$.
We constructed $\chi_1$ by
\begin{verbatim}
car25modI1(n)={
   my(a,b); a=n%25; b=0;
   if(gcd(a,5)>1,return(0),  
      while(2^b%25!=a,
         b++;
      );
   );
   return(8^b);
}

car7modI1(n)={
   my(a,b); a=n%7; b=0;
   if(gcd(a,7)>1,return(0),  
      while(3^b%7!=a,
         b++;
      );
   );
   return(47^b);
}

car175modI1(n)={
   return(car7modI1(n)*car25modI1(n));
}
\end{verbatim}
and we checked that $m_{\chi_1}\in I$ through

\begin{verbatim}
mcar175modI1={
   my(a);
   a=sum(x=1,175,x*car175modI1(x));
   a=a%61;
   return(a)
}
\end{verbatim}

To compute the various entries of the table we made use of
\begin{verbatim}
Acar175modI1(n)={
   my(a);
   a=sum(C=0,174,
        sum(D=0,174,
           car175modI1(D^2-C^2-n*C*D)*ceil((n*C-D)/175)*(C-175)
        );
     );
return(a%61);
}


Bcar175modI1(n)={
   my(a);
   a=sum(C=0,174,
        sum(D=0,174,
           car175modI1(D^2-C^2-n*C*D)*C*(C-175)
        );
     );
return(a%61);
}

Matrix1(v)={
   my(m,l); l=length(v);
   m=matrix(l,5);
   for(i=1,l,
      m[i,1]=v[i];
   );
   for(i=1,l,
      m[i,2]=Acar175modI1(2*v[i]);
   );
   for(i=1,l,
      m[i,3]=Bcar175modI1(2*v[i]);
   );
   for(i=1,l,
      m[i,4]=(v[i]-175*m[i,2]/(2*m[i,3]))%61;
   );
   for(i=1,l,
      m[i,5]=kronecker(v[i]^2+1,61);
   );
return(m);
}

\end{verbatim}

\end{document}